Circular No. 8695

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 F4 (SPACEWATCH) R. S. McMillan, Lunar and Planetary Laboratory, reports the Space-

watch discovery of a comet on CCD images taken by M. T. Read and himself with the 0.9-m reflector, the object being diffuse with a 6′′ tail in p.a. 280^{o} (discovery observation tabulated below). A. C. Gilmore reports that images taken on Mar. 30.6 UT with the Mt. John 1.0-m f/7.7 reflector show a moderately condensed coma ∼ 10′′ in diameter with a hint of a diffuse tail on the west side. Astrometry and preliminary orbital elements [T = 2006 Apr. 15.1 TT, ω = 25^{o}.3, Ω = 184^{o}.4, i = 12^{o}.5 (equinox 2000.0), q = 2.362 AU, e = 0.349, P = 6.9 yr] appear on MPEC 2006-F53.

2006 UT  α_{2000}  δ_{2000}  Mag.
Mar. 26.35589  14^{h}19^{m}57.^{s}50  −6^{o}55′40′′.7  19.9

RS OPHIUCHI R. Gonzalez-Riestra, XMM-Newton Science Operation Centre, Euro-

pean Space Agency; M. Orio, National Institute of Astrophysics of Italy and University of Wisconsin; and E. Leibowitz, Tel Aviv University, write that RS Oph was observed with XMM-Newton on Feb. 26 for 400 minutes and on Mar. 10 for 197 minutes. At both times, prominent emission lines were observed with the RGS gratings, due to H-like and He-like transitions of N, O, Mg, Ne, Si, and Fe. The total flux was ∼ 1.4 × 10^{−10} erg/s in the first observation and ∼ 1.1 × 10^{−10} erg/s in the second. Some emission lines in the range 0.6–1.6 nm decreased in strength by the second observation, while other lines in the region 1.7–3.3 nm became much more prominent, especially O VII (1.897 nm) and N VIII (2.478 nm), due to a decrease of the column density of neutral hydrogen to a value not exceeding 3 × 10^{21} cm^{−2} on Mar. 10. The lines were blueshifted by 1100–1400 km/s in the first observation and only by a few hundred km/s in the second, and at both times, their full-width-at-half-maximum corresponded to at least 2000 km/s. Both EPIC-pn spectra showed also the Fe-K-α complex in emission, but it was significantly stronger in the February spectrum. During the second observation, after the first 2800 s, the EPIC-pn count-rate in the range 0.15–0.50 keV quadrupled in ≈ 1700 s, then decreased by 40 percent in the following 1500 s and remained approximately constant until the end. During the rise and fall, the power spectrum shows a significant signal with a period of 35.685 s. The EPIC-pn count-rate in the range 1–10 keV had non-correlated variations by 10 percent and no coherent oscillations.

2006 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green